ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 18 Steiner, Richard, University of Glasgow, Great Britain Pasting in multiple categories A multiple category arises when several different categories have the same set of morphisms in a coherent way. I consider a particularly well-behaved type called an ω-category, which is the infinite-dimensional analogue of a strict n-category. These ω-categories can often be represented by abstract cell complexes called pasting diagrams, and in many cases they can be represented by genuine CW complexes. I characterise the cell complexes associated to ω-categories, and I give detailed descriptions of some particularly important examples. These examples elucidate the structure of tensor products.